A manufacturer of nickel-hydrogen batteries randomly selects 100 nickel plates for test cells, cycles them a specified number(b) If it is really the case that 16% of all plates blister under these circumstances and a sample size 100 is used, how likePart of the same question, need all of those parts answered. Double-check them as I have been getting the wrong answer. Thank you!
Answer #1

(x/n))=14/100 =0.14

z =(x/n-p)/√(p*(1-p)/n)= 1.33
p value = 0.0912

b)for n=100

rejection region: p+z*√p(1-p)/n = 0.1494
P(not be rejected given p=0.16)=P(phat<0.1494)=P(Z<(0.1494-0.1)/sqrt(0.16*0.84/100))= P(Z<-0.29)= 0.3859

for n=200

rejection region: p+z*√p(1-p)/n = 0.1349
P(not be rejected given p=0.16)=P(phat<0.1349)= P(Z<-0.97)= 0.1660

c)

required sample size             =n= ((zα(√po(1-po)+zβ(√pa(1-pa))/(p-po))2= 258